OpenStudy (anonymous):

Solve for x. 4(x-b)=x

OpenStudy (anonymous):

4x - 4b = x 3x = 4b x = 4/3b

OpenStudy (anonymous):

Thank you! I am just trying to see if I was right. I am grading my daughters paper

OpenStudy (anonymous):

4x-4b=x 4x-4x-4b=x-4x Subtract both sides by 4x to get -4b alone -4b=-3x Divide both sides by -3 to get "x" alone 4/3b = x Since two negative numbers cancelled each other out when multiply/divide the answer is positive.

OpenStudy (anonymous):

;)